# Probabilistic Neural Networks (PNNs) with t-Distributed Outputs: Adaptive Prediction Intervals Beyond Gaussian Assumptions

### v1

@@ Abstract
Traditional neural network regression models provide only point estimates, failing to capture predictive uncertainty. Probabilistic neural networks (PNNs) address this limitation by producing output distributions, enabling the construction of prediction intervals. However, the common assumption of Gaussian output distributions often results in overly wide intervals, particularly in the presence of outliers or deviations from normality. To enhance the adaptability of PNNs, we propose t-Distributed Neural Networks (TDistNNs), which generate t-distributed outputs, parameterized by location, scale, and degrees of freedom. The degrees of freedom parameter allows TDistNNs to model heavy-tailed predictive distributions, improving robustness to non-Gaussian data and enabling more adaptive uncertainty quantification. We develop a novel loss function tailored for the t-distribution and derive efficient gradient computations for seamless integration into deep learning frameworks. Empirical evaluations on synthetic and real-world data demonstrate that TDistNNs improve the balance between coverage and interval width. Notably, for identical architectures, TDistNNs consistently produce narrower prediction intervals than Gaussian-based PNNs while maintaining proper coverage. This work contributes a flexible framework for uncertainty estimation in neural networks tasked with regression, particularly suited to settings involving complex output distributions.
@@ 1 Introduction
In this work, we begin by outlining the necessary architectural modifications to implement TDistNNs, our proposed PNNs with t-distributed outputs. Specifically, we transform a deterministic neural network into a probabilistic one by augmenting the output layer. This augmentation involves extending the output from a single mean prediction to the prediction of three parameters: mean, scale, and degrees of freedom, which define the Student’s t-distribution. Each parameter is represented by a dedicated neuron in the output layer. We then derive a NLL function, specifically designed to optimize the parameters of the predictive t-distribution, thereby establishing a novel and principled approach for quantifying prediction errors within our probabilistic framework.
@@
In summary, we make the following contributions in this work:
  • This paper introduces a novel PNN framework designed to transform traditional deterministic neural networks into models capable of generating comprehensive predictive distributions. Specifically, we propose the utilization of the t-distribution as the foundational statistical model for these predictive distributions, enabling the simultaneous modeling of point predictions, their variability, and the output’s tail behavior. This approach extends and generalizes prior research that predominantly relied on Gaussian distributions, offering enhanced adaptability, particularly in the presence of outliers and heavy-tailed outputs.
  • Building upon the Student’s t-distribution framework, we proceed to derive the corresponding NLL loss function for model fitting. This derivation is followed by the analytical computation of the gradients of this loss function with respect to the parameters of the t-distribution, namely the mean, scale, and degrees of freedom. These analytically derived gradients are essential for enabling efficient training via backpropagation, and thus facilitate seamless integration with popular deep learning frameworks. This ensures that the developed TDistNN can be readily implemented and utilized within existing workflows.
  • To meticulously evaluate the performance of the proposed probabilistic framework, we conduct a series of comprehensive experiments. These experiments involve a detailed comparative analysis of our framework with existing methodologies, including those that employ parameterized distributions such as the Gaussian distribution. Furthermore, we extend our evaluation to include neural networks trained with the pinball loss function, which are commonly used for quantile regression. Through these experiments, we analyze the quality of the generated prediction intervals, with a specific focus on their widths and coverage levels. This analysis allows us to thoroughly investigate and present the inherent trade-offs between interval width and coverage accuracy, providing valuable insights into the practical applicability and limitations of the proposed framework and existing approaches.
The remainder of this paper is structured as follows. Section 2 reviews relevant notation, existing Gaussian-based PNNs, and pinball loss for quantile regression. Section 3 details our proposed model, which extends existing work by using the Student’s t-distribution as the predictive distribution, and provides the loss function derivation and gradient computations. In Section 4, we present numerical experiments that assess and compare the quality of prediction intervals generated by models with Gaussian and t-distributed outputs, as well as quantile regression. Finally, Section 5 summarizes the key findings, advantages, and future research directions.
@@ 2 Notation and Literature Review
Given the training data set $\mathcal{D}$, we estimate the parameters $\boldsymbol{\theta}$ in Eq. (2) by maximizing the likelihood of the observed data. The log-likelihood function for a single observation is [16]:
@@
Another (heuristic) approach to constructing prediction intervals is based on the pinball loss, which is commonly used in quantile regression. Instead of modeling the full distribution of the outputs, as done in Gaussian likelihood-based methods, quantile regression focuses on estimating conditional quantiles of the response variable or output given the input features [35]. For a given quantile level $\tau\in(0,1)$, the pinball loss for a single data point $(\mathbf{x}_{n},y_{n})$ is defined as [6]:
@@
For example, if we set $\tau=0.95$, the pinball loss penalizes underestimation much more heavily than overestimation. This is because when the prediction is too low, the residual is multiplied by $0.95$, making the penalty larger, whereas if the prediction is too high, the residual is multiplied by $1-0.95=0.05$, leading to a much smaller penalty. This behavior encourages the model to adjust predictions upward so that the estimated quantile represents a threshold below which most true values fall. By training separate models for different quantiles, such as $\tau=0.05$ and $\tau=0.95$, one can construct an empirical prediction interval that captures the response variable with an approximate 90% probability.
@@
In the final part of this section, we briefly discuss the integration of these loss functions, namely GaussianNLL and pinball loss, into traditional neural network architectures. For a regression task with $D$-dimensional input data and scalar outputs, a network comprises $D$ neurons in the input layer and a single neuron in the output layer. Multi-layered representations can be constructed by incorporating an arbitrary number of hidden layers with associated neurons between the input and output layers. The pinball loss function requires no architectural modifications to the neural network. However, it necessitates training separate models for each quantile level to estimate the lower and upper bounds of prediction intervals.
Conversely, the GaussianNLL loss function necessitates a modification to the network’s output layer. Specifically, the output layer must consist of two neurons: one dedicated to predicting the mean $f_{\mu}(\mathbf{x};\boldsymbol{\theta})$ and the other to predicting the variance $f_{\sigma}^{2}(\mathbf{x};\boldsymbol{\theta})$. These two outputs are then directly substituted into the GaussianNLL loss function, enabling the network to jointly optimize both the predictive mean and the associated uncertainty estimates within a single training process. This approach enables efficient, simultaneous learning of prediction intervals across multiple confidence levels. Fig. 2 illustrates the integration of these loss functions within neural networks.
Figure 2: We compare the architectures and methods for generating 90% confidence level prediction intervals using: (a) Quantile neural networks (QuantileNN), and (b) Gaussian-based neural networks (GaussianNN). QuantileNN employs a single output neuron to estimate the quantile at level $\tau$, optimized with the pinball loss. GaussianNN, in contrast, directly predicts the mean and variance through two output neurons, trained with the Gaussian negative log-likelihood (GaussianNLL) loss. The lower and upper bounds of the resulting prediction intervals are shown.
@@ 3 Probabilistic Neural Networks with t-Distributed Outputs
Specifically, we begin by discussing the parameterization of the t-distribution within the proposed PNN framework, detailing how the distribution’s location, scale, and shape parameters are derived from the network’s outputs. We then address the necessary modifications to standard deterministic neural network architectures to produce the designated predictive t-distribution. Subsequently, we present a detailed derivation of the negative log-likelihood (NLL) loss function, tailored for the t-distribution, and its gradients, which are essential for training the network via backpropagation. We conclude by detailing how to construct $(1-\alpha)\times 100\%$ confidence level prediction intervals using the learned t-distribution, for any $\alpha$ (e.g., $\alpha=0.1$), thus quantifying prediction uncertainty.
@@
To keep the notation concise in subsequent derivations, we will write $f_{\mu}$, $f_{\sigma}$, and $f_{\nu}$ without explicitly showing their dependence on the input feature vector $\mathbf{x}$ and the model parameters $\boldsymbol{\theta}$. Nonetheless, it is important to note that these quantities are indeed functions of both the input features and the learnable parameters of the neural network (weights and biases).
@@
[…]
@@
6:$\alpha\in(0,1)$: error rate for constructing prediction intervals (if desired)
@@
8:$\boldsymbol{\theta}^{\ast}$: optimal network weights/biases
@@
23:     Backpropagate: $\frac{\partial\ell_{\text{tDistNLL}}}{\partial\boldsymbol{\theta}}\;\leftarrow\;\text{via chain rule w.r.t.\ }f_{\mu},f_{\sigma},f_{\nu}$triangleright Equations (15), (17), (20)
24:     Update parameters: $\boldsymbol{\theta}\;\leftarrow\;\boldsymbol{\theta}\;-\;\eta\;\frac{\partial\ell_{\text{tDistNLL}}}{\partial\boldsymbol{\theta}}$triangleright Basic update formula
@@
28:$\bigl{(}f_{\mu},f_{\sigma},f_{\nu}\bigr{)}=\text{TDistNN}\bigl{(}\mathbf{x}_{\mathrm{test}};\boldsymbol{\theta}^{\ast}\bigr{)}$.
@@ 4 Experimental Results
In this section, we evaluate the performance of TDistNNs by systematically comparing it against existing methods. Our focus is on the construction of prediction intervals, a fundamental approach to quantifying uncertainty in regression tasks. Unlike point estimates, prediction intervals offer a probabilistic measure of confidence, enhancing model interpretability and ensuring more reliable decision-making in real-world applications where uncertainty is inherent, such as solving design optimization problems [26]. As detailed in the previous section, TDistNNs generate a point estimate via their first output node (see Fig. 3), while the scale and shape parameters of the fitted t-distribution define the interval bounds. This structure facilitates a direct and rigorous comparison with both Gaussian-based PNNs (GaussianNN) and quantile regression models optimized with the pinball loss (QuantileNN).
@@
A key objective of this paper is to show that integrating TDistNN into an existing neural network architecture is both straightforward and comparable to integrating other methods, such as QuantileNN and GaussianNN. To ensure a fair comparison, we minimize specialized hyperparameter tuning and maintain a consistent architecture throughout our analysis. In particular, we use a single-hidden-layer neural network throughout the experiments in this section.
@@
Unless stated otherwise, all models are trained using the Adam optimizer with a learning rate of 0.01 for 1,000 epochs, ReLU activation in the hidden layer, and a batch size corresponding to the entire training data set. Each method employs output-layer activations tailored to its respective distribution. For TDistNNs, we use the exponential function for the scale parameter and softplus for the shape parameter, ensuring valid t-distribution parameters. Similarly, in GaussianNNs, the exponential function is used to predict the variance of the output normal distribution. For QuantileNNs, we apply a standard linear activation function in the output layer. However, as mentioned, QuantileNNs require two separate training runs to estimate the desired prediction interval, corresponding to the quantile levels $\tau_{1}=0.05$ and $\tau_{2}=0.95$.
@@
Figure 4: Using a synthetic data set with outliers, we compare the performance of our proposed TDistNN against QuantileNN and GaussianNN. (a) shows the constructed prediction intervals, and (b) displays the histogram plot of degrees of freedom $f_{\nu}$ of the fitted t-distribution.
We observe that the QuantileNN model produces the narrowest prediction intervals throughout the entire input range while maintaining the target 90% coverage (recall that we set $\alpha=0.1$). In contrast, both GaussianNN and TDistNN models achieve slightly higher coverage scores, with 95% of true outputs falling within their respective prediction intervals. However, a significant advantage of the proposed TDistNN model is its reduced average interval width. At 5.20, it is 19.25% narrower than GaussianNN’s 6.44, while maintaining comparable coverage. This improvement is particularly noticeable at the input range’s extremities, where GaussianNN’s intervals widen to accommodate the outliers. The experimental results validate the t-distribution’s effectiveness in generating adaptive prediction intervals, outperforming the Gaussian model and providing a slight coverage enhancement over the QuantileNN model. Moreover, QuantileNNs necessitate multiple training iterations to approximate the predictive distribution due to its reliance on pre-defined quantiles.
@@
The coverage scores presented in Fig. 5(a) reveal that GaussianNN and TDistNN consistently surpass the 90% target coverage. While QuantileNN’s median coverage is adequate, its minimum values fall below the target coverage across all network architectures. Moreover, a consistent trend emerges: GaussianNNs exhibit higher coverage scores than TDistNNs. To evaluate whether this increased coverage is attributable to wider intervals designed to accommodate outliers within this data set, we must analyze the interval widths depicted in Fig. 5(b).
Figure 5: Boxplots illustrate (a) coverage score and (b) average interval width for different architectures across 20 trials on the synthetic data set with outliers. TDistNNs achieve the target coverage level of 90% with noticeably narrower prediction intervals than GaussianNNs, attributed to its modeling of heavier-tailed distributions.
The expected trend of GaussianNNs producing significantly wider prediction intervals than TDistNNs is observed, indicating that its higher coverage is a result of inflated uncertainty estimates. TDistNNs, in contrast, effectively construct more adaptive prediction intervals, striking a favorable balance between uncertainty quantification and interval tightness. Furthermore, a reduction in the variability of interval widths across the 20 independent trials is evident when employing 32 neurons in the hidden layer. In this architectural configuration, both QuantileNN and TDistNN achieve a median interval width of approximately 5, but TDistNN’s minimum coverage of 92% demonstrates a clear advantage in reliability. These results collectively highlight the merits of TDistNNs in complex data settings and elucidate the influence of network architecture on the key evaluation metrics.

### v2

# arXiv:2503.12354v2[cs.LG] 26 Feb 2026

@@ Abstract
Traditional neural network regression models provide only point estimates, failing to capture predictive uncertainty. Probabilistic neural networks (PNNs) address this limitation by producing output distributions, enabling the construction of prediction intervals. However, the common assumption of Gaussian output distributions often results in overly wide intervals, particularly in the presence of outliers or deviations from normality. To enhance the adaptability of PNNs, we propose t-Distributed Neural Networks (TDistNNs), which generate t-distributed outputs, parameterized by location, scale, and degrees of freedom. The degrees of freedom parameter allows TDistNNs to model heavy-tailed predictive distributions, improving robustness to non-Gaussian data and enabling more adaptive uncertainty quantification. We incorporate a likelihood based on the t-distribution into neural network training and derive efficient gradient computations for seamless integration into deep learning frameworks. Empirical evaluations on synthetic and real-world data demonstrate that TDistNNs improve the balance between coverage and interval width. Notably, for identical architectures, TDistNNs consistently produce narrower prediction intervals than Gaussian-based PNNs while maintaining proper coverage. This work contributes a flexible framework for uncertainty estimation in neural networks tasked with regression, particularly suited to settings involving complex output distributions.
@@ 1 Introduction
In this work, we begin by outlining the necessary architectural modifications to implement TDistNNs, our proposed PNNs with t-distributed outputs. Specifically, we transform a deterministic neural network into a probabilistic one by augmenting the output layer. This augmentation involves extending the output from a single mean prediction to the prediction of three parameters: mean, scale, and degrees of freedom, which define the Student’s t-distribution. Each parameter is represented by a dedicated neuron in the output layer. We then derive a NLL function, specifically designed to optimize the parameters of the predictive t-distribution.
@@
In summary, we make the following contributions in this work.
  • This paper introduces a novel PNN framework designed to transform traditional deterministic neural networks into models capable of generating comprehensive predictive distributions. Specifically, we propose the utilization of the t-distribution as the foundational statistical model for these predictive distributions, enabling the simultaneous modeling of point predictions, their variability, and the output’s tail behavior. This approach extends and generalizes prior research that predominantly relied on Gaussian distributions, offering enhanced adaptability.
  • Building upon the Student’s t-distribution framework, we proceed to derive the corresponding NLL loss function for model fitting. This derivation is followed by the analytical computation of the gradients of this loss function with respect to the parameters of the t-distribution, namely the mean, scale, and degrees of freedom. These analytically derived gradients are essential for enabling efficient training via backpropagation, and thus facilitate seamless integration with deep learning frameworks.
  • To evaluate the performance of the proposed probabilistic framework, we conduct a series of comprehensive experiments involving a detailed comparative analysis with existing methodologies. These include approaches based on parameterized output distributions, such as Gaussian probabilistic neural networks, neural networks trained with the pinball loss for quantile regression, as well as Monte Carlo Dropout-based uncertainty estimation methods. Through these experiments, we analyze the quality of the generated prediction intervals, with a specific focus on their widths and coverage levels. This analysis allows us to thoroughly investigate and present the inherent trade-offs between interval width and coverage accuracy, providing valuable insights into the practical applicability and limitations of the proposed framework and existing approaches.
The remainder of this paper is structured as follows. Section 2 reviews relevant notation, existing Gaussian-based PNNs, pinball loss for quantile regression, and Monte Carlo Dropout. Section 3 details our proposed model, which extends existing work by using the Student’s t-distribution as the predictive distribution, and provides the loss function derivation and gradient computations. In Section 4, we present numerical experiments that assess and compare the quality of prediction intervals. Finally, Section 5 summarizes the key findings, advantages, and future research directions.
@@ 2.1 Gaussian Output Distribution
Given the training set, we estimate the parameters $\boldsymbol{\theta}$ in Eq. (2) by maximizing the likelihood of the observed data. The log-likelihood function for a single observation is [18]:
@@
2.2 Quantile Regression
Another approach to constructing prediction intervals is based on the pinball loss, which is commonly used in quantile regression. Instead of modeling the full distribution of the outputs, as done in Gaussian likelihood-based methods, quantile regression focuses on estimating conditional quantiles of the response variable or output given the input features [37]. For a given quantile level $\tau\in(0,1)$, the pinball loss for a single data point $(\mathbf{x}_{n},y_{n})$ is defined as [5]:
@@ 2.2 Quantile Regression
For example, if we set $\tau=0.95$, the pinball loss penalizes underestimation much more heavily than overestimation. This is because when the prediction is too low, the residual is multiplied by $0.95$, making the penalty larger, whereas if the prediction is too high, the residual is multiplied by 10.95=0.051-0.95=0.05, leading to a much smaller penalty. This behavior encourages the model to adjust predictions upward so that the estimated quantile represents a threshold below which most true values fall. By training separate models for different quantiles, such as $\tau=0.05$ and $\tau=0.95$, one can construct an empirical prediction interval that captures the response variable with an approximate 90% probability.
@@
2.3 Monte Carlo Dropout
Monte Carlo Dropout (MC Dropout) is a widely used technique for uncertainty quantification in neural networks that does not require an explicit probabilistic output model [10, 12]. Originally introduced as a regularization strategy during training, dropout operates by deactivating a subset of neurons to reduce overfitting. This mechanism is controlled by a user-specified dropout rate, which determines the probability with which each neuron is deactivated during a forward pass through the network.
At each training iteration, neurons are stochastically omitted according to this Bernoulli process, while the remaining active units are appropriately rescaled to preserve the expected magnitude of activations. This stochastic architecture perturbation discourages co-adaptation of neurons and promotes the learning of robust, distributed feature representations. As a result, dropout can be viewed as implicitly training an ensemble of subnetworks that share parameters, with each subnetwork corresponding to a particular realization of the dropout mask.
MC Dropout extends this framework to the inference stage by retaining the same stochastic deactivation mechanism during prediction [9]. Rather than producing a single deterministic output for a given input, the model is evaluated multiple times with different dropout realizations, yielding stochastic predictions even for the same input. This induced variability enables the empirical approximation of a predictive distribution through repeated forward passes.
To be more formal, consider a neural network trained using the MSE loss in Eq. (1), augmented with dropout layers. At test time, dropout remains active and the model is evaluated $T$ times for a given input $\mathbf{x}_{\text{test}}$, producing a collection of stochastic predictions $\hat{y}^{(1)},\hat{y}^{(2)},\ldots,\hat{y}^{(T)}$, where $T$ denotes the number of Monte Carlo samples.
These samples define an empirical predictive distribution for the output conditioned on the input $\mathbf{x}_{\text{test}}$. Unlike likelihood-based regression methods, MC Dropout does not explicitly parameterize a conditional distribution such as a Gaussian. Instead, uncertainty is characterized directly from the variability of the Monte Carlo predictions. Consequently, prediction intervals are most naturally constructed using empirical quantiles of the sampled outputs, avoiding additional distributional assumptions. Specifically, a $(1-\alpha)\times 100\%$ prediction interval is given by:
$\left[Q_{\alpha/2}\bigl(\{\hat{y}^{(t)}\}_{t=1}^{T}\bigr),\;Q_{1-\alpha/2}\bigl(\{\hat{y}^{(t)}\}_{t=1}^{T}\bigr)\right],$ (7)
where $Q_{\tau}(\cdot)$ denotes the empirical $\tau$-quantile of the sampled predictions. This construction naturally accommodates asymmetric or heavy-tailed predictive behavior and is well suited when the shape of the output distribution is unknown.
In summary, MC Dropout provides a practical approach for uncertainty quantification in regression models trained with standard pointwise loss functions. While it enables flexible, assumption-free prediction intervals via sampling, the resulting uncertainty estimates are sensitive to the choice of dropout rate, network architecture, and number of Monte Carlo samples. Moreover, the absence of a likelihood-based training objective limits the principled calibration of predictive uncertainty.
@@ 3 Probabilistic Neural Networks with t-Distributed Outputs
Specifically, we begin by discussing the parameterization of the t-distribution within the PNN framework, detailing how the distribution’s location, scale, and shape parameters are derived from the network’s outputs. We then address the necessary modifications to standard deterministic neural network architectures to produce the designated predictive t-distribution. Subsequently, we present a detailed derivation of the negative log-likelihood (NLL) loss function, tailored for the t-distribution, and its gradients, which are essential for training the network via backpropagation.
@@
To keep the notation concise in subsequent derivations, we will write $f_{\mu}$, $f_{\sigma}$, and $f_{\nu}$ without explicitly showing their dependence on the input feature vector $\mathbf{x}$ and the model parameters $\boldsymbol{\theta}$. Nonetheless, it is important to note that these quantities are functions of both the input features and the learnable parameters of the network.
@@
The first derivative is simply $1/f_{\sigma}$. For the second, note that $\tfrac{\partial s_{n}}{\partial f_{\sigma}}=\tfrac{\partial}{\partial f_{\sigma}}\Bigl(1+\tfrac{z_{n}^{2}}{f_{\nu}}\Bigr)=\tfrac{1}{f_{\nu}}\,2\,z_{n}\,\tfrac{\partial z_{n}}{\partial f_{\sigma}}$. Since $z_{n}=\tfrac{r_{n}}{f_{\sigma}}$, we have $\tfrac{\partial z_{n}}{\partial f_{\sigma}}=\tfrac{\partial}{\partial f_{\sigma}}\Bigl(\tfrac{r_{n}}{f_{\sigma}}\Bigr)=-\,\tfrac{r_{n}}{f_{\sigma}^{2}}$. Putting it all together:
@@
[…]
@@
[…]
@@
23:  Backpropagate: $\frac{\partial\ell_{\text{tDistNLL}}}{\partial\boldsymbol{\theta}}\;\leftarrow\;\text{via chain rule w.r.t.\ }f_{\mu},f_{\sigma},f_{\nu}$ $\triangleright$ Equations (16), (18), (21)
24:  Update parameters: $\boldsymbol{\theta}\;\leftarrow\;\boldsymbol{\theta}\;-\;\eta\;\frac{\partial\ell_{\text{tDistNLL}}}{\partial\boldsymbol{\theta}}$ $\triangleright$ Basic update formula
@@
[…]
@@ 4 Experimental Results
In this section, we evaluate the performance of TDistNNs through a systematic comparison with existing methods. Our focus is on the construction of prediction intervals, a fundamental approach to quantifying uncertainty in regression tasks. Unlike point estimates, prediction intervals offer a probabilistic measure of confidence, enhancing model interpretability and ensuring more reliable decision-making in real-world applications where uncertainty is inherent, such as solving design optimization problems [28]. As detailed in the previous section, TDistNNs generate a point estimate via their first output node (see Fig. 2), while the scale and shape parameters of the fitted t-distribution define the interval bounds. This structure facilitates a direct and rigorous comparison with Gaussian-based PNNs (GaussianNN), quantile regression models optimized using the pinball loss (QuantileNN), and Monte Carlo Dropout (MC Dropout).
@@
A key objective of this paper is to show that integrating TDistNN into an existing neural network architecture is both straightforward and comparable to integrating other methods. To ensure a fair comparison, we minimize specialized hyperparameter tuning and maintain a consistent architecture throughout our analysis. Accordingly, we use a single-hidden-layer neural network for all experiments in this section, except for the final experiment, which considers deeper architectures.
@@
Unless stated otherwise, all models are trained using the Adam optimizer with a learning rate of 0.01 for 1,000 epochs, ReLU activation in the hidden layer, and a batch size corresponding to the entire training data set. Each method employs output-layer activations tailored to its respective distribution. For TDistNNs, we use the exponential function for the scale parameter and softplus for the shape parameter, ensuring valid t-distribution parameters. Similarly, in GaussianNNs, the exponential function is used to predict the variance of the output normal distribution. For QuantileNNs, we apply a standard linear activation function in the output layer. However, as mentioned, QuantileNNs require two separate training runs to estimate the desired prediction interval, corresponding to the quantile levels $\tau_{1}=0.05$ and $\tau_{2}=0.95$. For MC Dropout, we use a dropout rate of 0.2 with an ensemble size of $T=100$.
4.1 Synthetic Data with Heteroscedastic Noise and Outliers
@@
Figure 3: Using a synthetic data set with outliers, we compare the performance of our proposed TDistNN against QuantileNN, GaussianNN, and MC Dropout. (a) shows the constructed prediction intervals, and (b) displays the histogram plot of degrees of freedom $f_{\nu}$ of the fitted t-distribution.
We observe that the QuantileNN model produces the narrowest prediction intervals throughout the entire input range while maintaining the target 90% coverage (recall that we set $\alpha=0.1$). While MC Dropout produces narrower intervals, this comes at the cost of substantial undercoverage (78% versus the target 90%), and its resulting prediction bounds are noticeably less smooth than those of the other methods. In contrast, both GaussianNN and TDistNN models achieve slightly higher coverage scores, with 95% of true outputs falling within their respective prediction intervals. However, a significant advantage of the proposed TDistNN model is its reduced average interval width. At 5.20, it is 18.24% narrower than GaussianNN’s 6.36, while maintaining comparable coverage. This improvement is particularly noticeable at the input range’s extremities, where GaussianNN’s intervals widen to accommodate the outliers.
Therefore, the experimental results validate the t-distribution’s effectiveness in generating adaptive prediction intervals, outperforming the Gaussian model and providing a slight coverage enhancement over the QuantileNN model. Moreover, QuantileNNs necessitate multiple training iterations to approximate the predictive distribution due to its reliance on pre-defined quantiles.
@@
The coverage scores presented in Fig. 4(a) reveal that GaussianNN and TDistNN consistently surpass the 90% target coverage. In contrast, MC Dropout fails to reach the 90% target coverage in any of the 20 trials. While QuantileNN’s median coverage is adequate, its minimum values fall below the target coverage across all network architectures. Moreover, a consistent trend emerges: GaussianNNs exhibit higher coverage scores than TDistNNs. To evaluate whether this increased coverage is attributable to wider intervals designed to accommodate outliers within this data set, we must analyze the interval widths depicted in Fig. 4(b).
Figure 4: Boxplots illustrate (a) coverage score and (b) average interval width for different architectures across 20 trials on the synthetic data set with outliers. For each network size, the results are shown in the following left-to-right order: QuantileNN, GaussianNN, MC Dropout, and TDistNN.
The expected trend of GaussianNNs producing significantly wider prediction intervals than TDistNNs is observed, indicating that its higher coverage is a result of inflated uncertainty estimates. TDistNNs, in contrast, effectively construct more adaptive prediction intervals, striking a favorable balance between uncertainty quantification and interval tightness. When employing 32 neurons, both QuantileNN and TDistNN achieve a median interval width of approximately 5, but TDistNN’s minimum coverage of 92% demonstrates a clear advantage in reliability. These results collectively highlight the merits of TDistNNs in complex data settings and elucidate the influence of network architecture on the key evaluation metrics. In addition, we observe that for MC Dropout, when the number of neurons is 32, the variability across the 20 trials in both coverage and interval width is higher than that of the other models.
4.2 UCI Regression Benchmarks